OpenStudy (anonymous):

you should always multiply before you divide , is this statement true or false ? i think its true you should always multiply before you divide , is this statement true or false ? i think its true @Mathematics

OpenStudy (anonymous):

actually, it doesnt matter if you multiply then divide, or the other way around. "Division" is just "Multiplication by inverses". Example: 3 and 1/3 are inverses because:\[3\cdot \frac{1}{3}=1\] So I can either do \[90\div3=30\]or\[90\cdot \frac{1}{3}=30\] So if i have a long string of multiplication and division:\[100\cdot5\cdot10\cdot8\div5\div10\div8\]i can do it in any order:\[100\cdot5\div5\cdot10\div10\cdot8\div8=100\cdot1\cdot1\cdot1=100\]

OpenStudy (anonymous):

BODMAS rule?